Find the centre and radius of the equation of the circle. 2x^2+2y^2-8x+5y+10=0
1
Expert's answer
2020-12-10T14:13:37-0500

Solution. Transform the equations of the circle


"2x^2+2y^2-8x+5y+10=0""2(x-2)^2-8+2(y+\frac{5}{4})^2-\frac{25}{8}+10=0""2(x-2)^2+2(y+\frac{5}{4})^2=\frac{9}{8}""(x-2)^2+(y+\frac{5}{4})^2=\frac{9}{16}"

General equation of a circle


"(x-a)^2+(y-b)^2=R^2"

where (a,b) is circle center coordinates; R is circle radius.

Therefore circle center coordinates


"(2,-\frac{5}{4})"

circle radius


"R=\frac{3}{4}"

Answer.

"(2,-\frac{5}{4})""R=\frac{3}{4}"